how many liters of 20% alcohol soltuion must be mixed with 10 liters of a 50% solution to get a 30% solution?
Answer by Alan3354(69443)

how many liters of 20% alcohol soltuion must be mixed with 10 liters of a 50% solution to get a 30% solution?
-------------------------------
0.20*x + 0.50*10 = 0.30*(x + 10)
2x + 50 = 3x + 30
x = 20 liters of 20%